ocean seismicity banner imageVents ProgramNOAAOcean SeismicityAcoustics
  Deep Ocean Seismicity from Hydroacoustic Monitoring
background image background image background image background image background image background image background image background image background image
background image background image
background image go to Acoustics home background image scroll down for Methods background image scroll down for Results background image scroll down for Multimedia background image spacer image
background image background image background image background image spacer image
  Response Cruise - Endeavour Ridge- February-March, 2005  
cruise summary map
Response cruise summary map:
white stars = CTD casts
white lines & stars = CTD tows
blue lines & dots = camera tow
red dots = earthquake epicenters
Bathymetry is NOAA multibeam.

Back to main event page: Endeavour, March 2005
Visit: Endeavour Science Update site

News from sea, Jim Cowen (Chief Scientist)
Cruise TN-177B, R/V Thompson
March 2005 Endeavour Rapid Response Cruise:

March 9:
~0044 hr: End V05B-02.
~0144 hr: Start third off-axis vertical cast VO5B-03: 47o 58.4’ N; 129o 18.5’ W
~0311 hr: End cast V05B-03. None of the last three casts (all verticals) showed any real-time (CTD-optical) evidence of a particle or thermal plume, which would be indicative of hydrothermal plumes that had been carried off-axis by prevailing ocean currents. Chemical analyses of the water samples collected during these casts will take some time.

We then moved operations back to the main Endeavour axial valley, with intention of performing a series of vertical casts over most of the known active vent fields.
~0500 hr: Started vertical cast V05B-04: 47o 56.866’ N; 129o 05.962’ W, over the Main Endeavour field.
~0630 hr: End V05B-04. A vigorous thermal and particle plume was found, typical of this site. Further analyses are needed to determine whether the recent earthquake swarm caused a change in the subseafloor’s hydrothermal plumbing, in terms of heat flux and chemical composition.
~0936 hr: Start V05B-05. Vertical cast over the Mothra vent field, about 2.5 km south of the MEF (47o 55.4140’ N, 129o 06.5310’ W).
~1115 hr: End V05B-05. Again, like the MEF, this cast over Mothra found a vigorous hydrothermal plume. The plume was less intense than at MEF in terms of the particle anomaly, but from the strong presence of hydrogen sulfide (that rotten egg smell) in some of the water samples collected, it appears that we were able to obtain samples of the buoyant phase or very young neutrally buoyant phase of the plume. This should facilitate comparisons of values for current hydrothermal fluid compositional ratios with historic values. However, ship-board analysis of methane and hydrogen indicate that vent fluid chemistry at MEF and Mothra do not significantly differ from “normal”.

camera photo camera photo
Images from camera tow. Camera photos show moderate to heavily sedimented pillows and talus throughout the whole traverse. Lots of sessile biota (sponges, corals) and even viewed a small octopus in one photo. No evidence of recent lavas or venting in the area we traversed. (click images for full view)

~1630 hrs: Deep-sea camera is launched. CTD-rosette casts have been temporarily halted in order to stage and launch the deep-sea camera. While the ‘water column’ crew repeatedly launched, recovered, and processed CTD-rosette casts through the night and morning hours, the camera crew troubleshooted the problem with the camera system. After much work and creative solutions, the deep-sea camera was launched for an 8 hour tow along the seafloor from 48o 05.628’ N, 129o 01.596’ W toward 48o 01.7, 129o 03.5’ W. This path is north of the known Endeavour vent fields and directly over a suspected magma chamber. Oops, back up for a little more problem solving.
~1930 hrs: Camera back in the water. This time everything was working, including the CTD. During this tow several small thermal anomalies may have been picked up by the camera system’s CTD. All indications are that the digital camera and strobes themselves are functioning properly.

image of camera sled
Camera system aboard the response cruise.

~0430 hrs: Camera system back on deck after successful 8 hour tow. The data collected during the camera tow are consistent with the other cruise results. Camera images of the seafloor along its single tow path showed only sediment covered seafloor, no fresh basalts. Neither the camera sled’s CTD data nor data from the Miniature Autonomous Plume Recorder (MAPR) attached to the camera sled showed any significant thermal or optical anomalies (the MAPR was mounted on the sled in case the CTD malfunctioned during the tow).
Following the recovery of the camera system, the vertical CTD-rosette casts were continued over the northern known Endeavour vent fields:

~0530 to ~0700 hrs: Vertical cast V05B-06 (High Rise vent field at 47° 58.1259 N, 129° 05.2027).
~0910 to ~1100 hrs: Vertical cast V05B-07 (Salty Dawg vent field at 47° 58.922 N, 129° 04.569).

Both of these casts intercepted buoyant hydrothermal plumes, thus supplying concentrated plume samples for chemical analyses.

Bottom line as we head for shore:
1. It appears unlikely that this February/March 2005 earthquake swarm induced corresponding expression at the seafloor (e.g., eruptive flow) or in the water column (e.g., hydrothermal chronic or event plume). The in situ and ship-board physical and chemical data from the 3 long tow-yo casts and 7 vertical casts revealed no water column signal that can be clearly associated with the recent earthquake swarm, whether magmatic or tectonic. Initial calculation of methane to hydrogen ratios from MEF and Mothra are comparable to historic values from vent fluids. On the other hand, methane to hydrogen ratios from CTD-bottle casts at High Rise and Salty Dawg may differ from pre-event vent fluid ratios, which could imply a connection to the seismic event, especially since these vent fields are farther north than MEF and Mothra (and thus closer to the event earthquake epicenters). No evidence of any temperature or optical anomalies were seen in the near-bottom camera tow data (CTD or MAPR). Camera images of the seafloor revealed no fresh basalt; rather, the entire camera tow track was sediment covered. Finally, we searched for evidence of new lava flows in the earthquake area by comparing high-resolution multibeam bathymetry data with the historic Seabeam data. No bathymetric anomalies were detected.
Time-series records from in situ sensors currently deployed at various vents along the Endeavour segment may have recorded event-related fluctuations in temperature (or other parameters); all investigators with sensors deployed at Endeavour should check these records carefully once they are recovered.

2. Extensive shore-based chemical analyses of samples collected during this short but intense cruise, as well as further scrutiny of CTD, optical, and ship-board chemical analyses, are needed before the possibility of event-related venting can be finally confirmed or eliminated.

3. Possible explanations for failure to discover significant event-related hydrothermal discharge include:
• The earthquake swarm was likely an intrusive, magmatic event that may not have 1) reached sufficiently shallow crustal depths to lead to extrusion (eruptive flows), 2) caused changes in hydrothermal circulation at the seafloor and/or 3) did not stimulate new venting or changes to existing venting as discernable via surface ship sampling.
• Hydrothermal changes associated with this event were focused at the 5 known Endeavour Vent fields. Detailed sampling of the plumes overlying the known Endeavour vent fields was performed following an earthquake swarm west of Endeavour in 1999 demonstrating clear changes in the fluid chemistry of several of these vents following that 1999 swarm. It may be that even subtle changes in the ratios of key hydrothermal species following the earthquakes will be detectable in these plume samples since there exists an extensive history of the chemistry of these fluids and plumes. However, our preliminary ship-board analyses clearly do not support this hypothesis.
• Inadequate sampling. Our rapid response was not fast enough and event-related venting had stopped in the week between the start of the earthquake swarm and our on-site arrival, so that any event plume was swept away out of reach of our time-limited search net. This option seems unlikely since the pattern of our tow-yos should have detected any sustained significant venting occurring within the region defined by the earthquake swarm, given the most likely current patterns and regional geology/bathymetry. Furthermore, event-related hydrothermal discharge has been sustained for months to years following the original seismic event for all of the few seafloor eruption events that have so far been documented.

4. This episode of event remote detection and rapid response demonstrated a remarkable degree of cooperation and dedication among university and government scientists and their respective infrastructures, funding agencies and their program managers, UNOLS and university ship operators, and a host of other important contributors to this effort. The Ridge2000 community was informed of the seismic swarm on Monday, February 28. Animated discussion concerning the nature and implications of this event started immediately. Response personnel were enroute by Thursday, loading the ship on Friday, and sailed from the University of Washington dock at 0900 hr Saturday morning. We were on station by Sunday morning, just 6 days after notification of the seismic swarm, a task that usually requires a lead-time of over a year. The detection and response team gratefully acknowledges the assistance of the very many who got us on our way. We want to especially thank Dave Epp of the NSF-Ridge2000 Program and Steve Hammond of NOAA-Ocean Exploration for

image of R/V Thompson
R/V Thompson at the dock.

funding the response cruise, Daniel Schwartz and the UW marine facility for logistical help of all kinds, the US State Department and the Canadian Ministry of Fisheries and Ocean for expediting the permit process (operation in foreign waters), UW’s and SIO’s Radiation Safety Offices, and of course Captain Phil Smith and the entire crew of the RV Thomas T Thompson for incredible hospitality and skilled assistance throughout this cruise even though we had so suddenly upended their home port time. Most of all we want to thank our families and close friends for their support during the temporary loss of normal life. Finally, the shipboard PIs would like to thank each and every one of the shipboard scientists who upended their own lives to enthusiastically join the cruise; you have all been lively and professional.

March 8
~1430 hrs: Ended cast T05B-03. Then deployed the deep-sea camera (beautiful deployment), but had to abort because of a problem with its CTD. Camera crew immediately set to work to troubleshoot problem.

~1930 hrs: It was subsequently determined that camera testing could take several additional hours. Decision was made to switch back to CTD-bottle casts while camera crew remains hard at troubleshooting, and hopefully repairs. A series of vertical casts will be made further to the Southwest of the second tow at:
V05B-01: 48o 7.9’ N; 129o 13.0’ W
V05B-02: 48o 3.0’ N; 129o 15.6’ W
VO5B-03: 47o 58.4’ N; 129o 18.5’ W
The motivation behind these vertical casts to expand the search net for any potential event plumes. These vertical casts will be followed by a series of casts or a tow along the main axis of the Endeavour over the known vent fields in order to test the hypothesis that the intense seismic activity north of these fields could have induced changes in the intensity and chemical character of the hydrothermal discharge here, as observed following the 1999 Endeavour seismic swarm.

Spirits and motivation remain high. Science crew has meshed very well and has become quite an efficient team. Ship’s crew is fantastic, demonstrating great cooperation and excellent skill—a can-do captain and crew. Weather has also improved, for the moment at least. Wonderful pumpkin pie last night!

March 7 continued
~1710 hr: Cast T05B-02 on deck. Full sampling started immediately as above, but also included analyses for radon, extremophile cultures, dissolved organic carbon, and scanning electron microscopy, as well as manganese-54 tracer experiments. The instrument for in situ particle size distributions appeared to work well during the cast. Again no clear hydrothermal signal. The bottom nephaloid layer intensified over the southern third of the tow. Ship-board analyses showed above background values for methane (4-14 nM) along the tow tract and at variable depths above bottom ((<50 to ~300 m), both with and without corresponding particle plumes. No clear nearby source for this plume is yet evident, although temperature-salinity plots, the spatial pattern of light scattering plumes, and the nature of the suspended particles suggest that at least some of the particle plumes (at ~ 2400 m) and elevated methane may derive from resuspension of sediments (and associated reduced gases—i.e., methane) on the broad shallow sill to the east of the earthquake swarm area.

~1720-1930 hrs: Ran a EM300 multibeam (high resolution bathymetry) track from .

~2000 hrs: Started a new tow-yo cast, T05B-03. Cast start position: 48o 12.50 N; 129o 3.47’ W, heading in a southerly direction to 48o 5.8’ N, 129o 4.6’ W then on to 47o 5.3’ N, 129o 9.3’ W. This tow bisects region between the first two, and then continues much further south. Following this tow, we anticipate a series of vertical tows both expanding our initial survey of the region as well as further investigating certain areas closer to the seafloor. Weather permitting we also anticipate performing a camera tow starting this afternoon. Currently the winds are out of the SSE at about 30 knots.

March 7
Cast T05B-01 was recovered and immediately subsampled for He, CH4/H2, CH4 stable isotopes, ammonia, CO2, pH, voltammetric electrochemistry, Total, dissolved, and particulate metals, microbial biomass and molecular biology, transmission electron microscopy. Initial ship-board analyses do not reveal any dramatic hydrothermal signals, but preliminary results suggest that one or more samples from (the 2400 m particle plume and the Eh spike) may have elevated methane and soluble FeS, elemental S. More calibrations need to be performed before such preliminary results are confirmed. EM300 bathymetry data is great.

~0300 to ~0430 hr: An EM300 (bathymetry) line was run from 48 05.00 N, 129 05.0700 north to 48 15.000 N, 129 00.4200 W, during cast turn around and transitting to new cast start postion.

~0502 hr: Cast T05B-02 started. Start point: 48 12.990 N, 129 07.487 W; End point: . This is a tow from start WP due south. A small, but significant above-bottom particle plume signal appeared about half way through this tow; no other electronic signals appeared to correspond to this particle plume, but several Niskin bottles were tripped for on-board sample analyses. This tow is still underway. Strong winds (gusts to ~50 knots) and rising seas (20-30 ft) are expected later this afternoon/evening; we will likely be restricted to vertical casts at that time. A lessening in the wind and sea conditions is expected in morning or early afternoon of following day.

Late March 6 2350 hrs: Arrived on station after some working through some first cast issues, we started the first cast (T05B-01). EM300 multibeam (high resolution bathymetry) data was collected during this and all tows. Once fish was near seafloor, we started the tow-yo, tripping 3 bottles per each ascent cycle. During this cast Marshall Schwartz (WHOI) was our early cruise hero; he was tireless and instrumental in troubleshooting and repairing an in situ electronic instrument of the Hawaii group, with the sustained help of UH?s Yves Plancherel and Brian Glazer, All sample (Niskin) bottles were tripped. A significant nephaloid layer was observed near bottom throughout cast. A secondary vertically narrow particle plume was also observed for about the first half of the tow at ~2400 m. There was also sharp change in Eh at one location (descent only)?may have been instrument problem because the baseline appeared to have readjusted itself at this time.

March 6: Continuing toward Endeavour Segment (Way Point 1: 48o 12’N, 128o 58.9’W). Estimated time of arrival is 1400 hrs (two hours from now). It is overcast, 9 degrees C, with ~12 ft seas and ~22 knot winds.
Weather permitting we will do a tow-yo (CTD-rosette) cast immediately upon arrival on station, towing from WP 1 toward the SSW, essentially over the ridge axis. I sea and wind conditions preclude a tow, we will commence a series of vertical CTD-rosette casts. Watches are set and all hands ready to go. Stay tuned.

March 5: Departed from UW dock at 0900 hr. Proceeded through locks to Straights of Juan de Fuca.

The following scientists are participating on the event response cruise:
Jim Cowen Chief Scientist/Geochem
Ed Baker co-Chief/Plume Mapping
Bill Chadwick Geology
Joe Haxel Tracer Geochemistry
Joe Resing Geochemistry
Sharon Walker Plume Mapping
Geoffrey Lebon Geochemistry
Kelley Kaiser Geochemistry
Shannon Ristau Seafloor Mapping
Brian Glazer Microbial Geochemistry
Yves Plancherel Biochemsitry
Phyllis Lam Microbial Geochemistry
Kristen Mailheau Microbial Geochemistry
Eric Olson Geochemistry
Deborah Glickson Geology
WesThompson Geochemistry
Mark Stephens  Geochemistry
Jenn Josef Microbiology
Marshall Swartz Seafloor imaging
Rhian Waller Seafloor imaging
Melissa Rotella Microbiology

Back to main event page: Endeavour, March 2005

Note there was another swarm of earthquakes in this same area in October 2004.

Last Updated: 03/03/05

Address inquiries to:
Robert Dziak - Seismologist - robert.p.dziak@noaa.gov
Matt Fowler - Analyst - matt.fowler@noaa.gov


Vents Home Contacts-Credits Bibliography Links Disclaimer Privacy Policy